# The Ginsparg-Wilson relation and overlap fermions

### v1

@@ IV Numerical methods for overlap fermions
The latter choice is easier to implement and interpret.
@@
The Zolotarev formula represents the state of the art. Referring readers to the mathematical literature (see  [Akhiezer1, Akhiezer2]) for its derivation, I simply present the recipe for its implementation. The coefficients are defined in terms of the elliptic integrals $\rm cel$}, $\rm cn$}, $\rm sn$}, and $\rm dn$}, which I will list due to the multitude of conventions (I am following those of Ref. [Press:1988xx]): The approximate step function is
@@
over the desired range, so it is straightforward to pick a desired accuracy for a known range simply by varying $N$. We illustrate an example of a Zolotarev approximation in Fig. (1).
@@
Evaluating the squared Hermitian overlap operator $H^{2}=(\gamma_{5}D)^{2}=D^{\dagger}D$ underpins essentially all numerical work with the overlap. Recall that $H^{2}$ commutes with $\gamma_{5}$ and its eigenvectors can be chosen with definite chirality. It is convenient to define the chiral projections ($P_{\pm}=\frac{1}{2}(1\pm\gamma_{5})$) so that the massive squared Hermitian overlap operator, with the usual convention for the mass terms, is $H(m)^{2}=H_{+}(m)^{2}+H_{-}(m)^{2}$ where
@@
Eigenmodes and eigenvalues of $H$ have a literature of their own. I have had good success with two methods to find eigenmodes and eigenvalues of $H$. One is an implementation of the Ritz variational method as described by the authors of Ref. [Kalkreuter:1995mm] and the other is the “primme” package of Refs. [doi:10.1137/050631574, doi:10.1137/060661910].
@@
Overlap fermions are particularly suited to a limit of QCD called the “epsilon regime” [Gasser:1986vb, Gasser:1987ah]. This is a situation involving a finite volume $V=L^{3}$ and a quark mass so light that $m_{\pi}L<<1$ while $ML>>1$ for all other hadrons. The partition function is dominated by the pions. Everything else is massive and their contribution to the partition function is exponentially suppressed. In that limit there is a connection between eigenvalues of the Dirac operator (typically in fixed sectors of topology) and random matrix theory. The first description of this connection was by Shuryak and Verbaarschot [Shuryak:1992pi] and there is a large subsequent literature. An old review by Damgaard [Damgaard:2001ep] is still a good introduction to the subject. The most recent discussion of this topic is probably the 2019 FLAG review [FlavourLatticeAveragingGroup:2019iem]. In a sentence, the probability distribution of Dirac eigenvalues can be used to determine the condensate and other low energy constants of the chiral Lagrangian.
The advantage of using the overlap as opposed to other formulations of lattice fermions is its exact chiral symmetry, which strongly constrains patterns in the spectrum – zero modes are chiral, nonzero eigenvalue modes are not. This spares a lot of argument about how lattice artifacts affect results. (For example, is a small eigenvalue eigenmode an approximation to a true zero mode, or not?) An issue with the program is that in the epsilon regime, finite size corrections are power law in $1/L$, rather than exponential ($\exp(-m_{\pi}L)$) for conventional measurements. In the end, large volumes were needed to overcome these large corrections.
@@
Now for a few words about simulations with dynamical overlap fermions. All the projects I know of used hybrid Monte Carlo with the $\Phi$ hybrid molecular dynamics (MD) algorithm. The $\Phi$ algorithm obtains the force on the gauge field from the derivative of the pseudofermion effective potential $S_{f}=\Phi^{*}[M(U)^{\dagger}M(U)]^{-1}\Phi$ with respect to the vector potential $A$ or $U$. Because the fermion matrix depends on the gauge field, the pseudofermion term contributes the “fermion force”,
@@
A reversible molecular dynamics treatment of a discontinuous effective potential was proposed by [Fodor:2003bh]. It is based on the analogy with classical particle motion in the presence of a step barrier. Imagine a unit-mass particle moving in two spatial dimensions $(x,y)$, and approaching a potential step, with $V(x,y)=0$, say, for $x<0$ and $V(x,y)=V_{0}$ for $x>0$. The component of momentum in the $y$ direction is unchanged when the particle strikes the barrier, but there is an impulse in the $x$ direction, which causes a discontinuous change in $p_{x}$. If the barrier is too high, the particle “reflects,” $p_{x}\rightarrow-p_{x}$, but if the barrier is low enough, the particle crosses the barrier (“refracts”) while $p_{x}$ changes. The change is given by energy conservation, $\frac{1}{2}p_{x}^{2}=\frac{1}{2}(p_{x}^{\prime})^{2}+V_{0}$.
@@
In practice, the exact step function of the overlap operator is replaced by some approximation, such as a rational functional approximation. To obtain an accurate approximation to the step function, it continues to be necessary to project out low modes of the kernel operator, as in Eq. (36). This introduces another complication for HMC: we have to be able to differentiate the projector. Evaluating the derivative requires determining the change in the projector under small perturbations in $h$. Labeling the projector on an eigenmode of h(r0)h(-r0) with eigenvalue $\lambda$ as $P_{\lambda}=|\lambda\rangle\langle\lambda|$ , first order perturbation theory gives [Narayanan:2000qx]
@@
All this business with reflection and refraction is morally equivalent to the problem of changing topology as seen in simulations with actions which treat topology only approximately (meaning, everything else except overlap fermions). What is different is that here, the barrier between topology changes is sharp and easy to understand. As with other actions, simulations with overlap fermions have a hard time changing topology and the problem gets worse as the lattice spacing decreases. It is very easy, though, to perform simulations in sectors of fixed topology, simply by always reflecting from topological boundaries. These data sets can used in the epsilon regime studies described above, to check predictions for the properties of eigenvalues in sectors of fixed winding number.
The magical properties of overlap fermions allow simulations of an odd number of flavors without rooting. For details, see Ref. [DeGrand:2006ws].
Finally, let me make a few remarks about my own experiences with the overlap as a computational problem. (This is described in more detail in Ref. [DeGrand:2000tf].) I was never able to get anything useful out of an overlap operator whose kernel $d$ or $h$ was a usual Wilson fermion with thin links. It was too expensive to evaluate; $\epsilon_{N}$ needed very large $N$ and very small $c_{2l-1}$’s. The issue, I always thought, was that while the spectrum of the overlap $D$ was a circle, that was not the case for the kernel Wilson action. An example is shown in Fig. 2. I always felt that it was asking a lot of the computer to map the crosses in Fig. 2 onto the Ginsparg-Wilson circle. Actions with nearest neighbor and diagonal ($\pm\hat{i}\pm\hat{j}$) couplings produced approximately circular spectra to start with, and it was easier for the overlap operator to circularize them. Smeared gauge connections plus a clover term further smoothed the action and allowed me to use an $\epsilon_{N}$ with small $N$ and a more well-behaved range of eigenvalues of $h$. I used eigenmodes of $h$ to precondition the calculation of $H$ and eigenmodes of $H$ to precondition the calculation of $\hat{D}^{-1}$.
@@ V Final thoughts
As far as I can tell from a literature search, only one lattice group doing large scale simulations used overlap fermions (though there were several groups doing small scale simulations). That was JLQCD, who started carrying out $N_{f}=2$ flavor simulations around 2006, with a $16^{3}\times 32$ volume and quark masses down to a nominal 20 MeV or $m_{q}=m_{strange}/6$ – see Ref. [JLQCD:2006lvu]. A year later they were doing $N_{f}=2+1$ flavor simulations. Their overlap program continued until about 2014, at which point they converted to simulating Moebius domain wall fermions [Noaki:2014sda]. These fermions do not have exact chiral symmetry, but the additive mass renormalization they observed was about 1/2 MeV, which they argued was more than adequate for their needs.
Advances in techniques also rendered many of the theoretical advantages of exact chiral symmetry less important. For example, these days the topological susceptibility can be measured using gradient flow [Luscher:2010iy] from simulations with nonchiral fermion actions, rather than by counting zero modes. The condensate can be measured from the integrated spectral density of eigenvalues of a (nonchiral) Dirac operator [Giusti:2008vb]. And of course, most of the “classic problems” from the time of the development of the overlap have been solved: measurements of the condensate or the pseudoscalar decay constant are not so newsworthy any more. Groups have moved on to more complicated observables, where many things have to be balanced against each other – small lattice spacing, large volume, large statistics, multiparticle correlation functions, good chiral symmetry properties, and more.

### v2

@@ IV Numerical methods for overlap fermions
This choice is easier to implement and interpret.
@@
The Zolotarev formula represents the state of the art. Referring readers to the mathematical literature (see  [2, 1]) for its derivation, I simply present the recipe for its implementation. The coefficients are defined in terms of the elliptic integrals $\rm cel$}, $\rm cn$}, $\rm sn$}, and $\rm dn$}, which I will list due to the multitude of conventions (I am following those of Ref. [30]): The approximate step function is
@@
over the desired range, so it is straightforward to pick a desired accuracy for a known range simply by varying $N$. We illustrate an example of a Zolotarev approximation in Fig. 2.
@@
Evaluating the squared Hermitian overlap operator $H^{2}=(\gamma_{5}D)^{2}=D^{\dagger}D$ underpins essentially all numerical work with the overlap. Recall that $H^{2}$ commutes with $\gamma_{5}$ and its eigenvectors can be chosen with definite chirality. It is convenient to make use of the chiral projections ($P_{\pm}=\frac{1}{2}(1\pm\gamma_{5})$) so that the massive squared Hermitian overlap operator, with the usual convention for the mass terms, is $H(m)^{2}=H_{+}(m)^{2}+H_{-}(m)^{2}$ where
@@
Eigenmodes and eigenvalues of $H$ have a literature of their own. I have had good success with two methods to find eigenmodes and eigenvalues of $H$. One is an implementation of the Ritz variational method as described by the authors of Ref. [15] and the other is the “primme” package of Refs. [33, 32].
@@
Overlap fermions are particularly suited to a limit of QCD called the “epsilon regime” [10, 11]. This is a situation involving a finite volume $V=L^{3}$ and a quark mass so light that $m_{\pi}L<<1$ while $ML>>1$ for all other hadrons. The partition function is dominated by the pions. Everything else is massive and their contribution to the partition function is exponentially suppressed. In that limit there is a connection between eigenvalues of the Dirac operator (typically in fixed sectors of topology) and random matrix theory. The first description of this connection was by Shuryak and Verbaarschot [31] and there is a large subsequent literature. An old review by Damgaard [5] is still a good introduction to the subject. The most recent discussion of this topic is probably the 2019 FLAG review [3]. In a sentence, the probability distribution of Dirac eigenvalues can be used to determine the condensate and other low energy constants of the chiral Lagrangian.
The advantage of using the overlap as opposed to other formulations of lattice fermions is its exact chiral symmetry, which strongly constrains patterns in the spectrum – zero modes are chiral, nonzero eigenvalue modes are not. This spares a lot of argument about how lattice artifacts affect results. (For example, is a small eigenvalue eigenmode an approximation to a true zero mode, or not?) An issue with the program is that in the epsilon regime, finite size corrections are power law in $1/L$, rather than exponential ($\exp(-m_{\pi}L)$) as they are for conventional measurements. In the end, large volumes were needed to overcome these large corrections.
@@
Now for a few words about simulations with dynamical overlap fermions. All the projects I know of used hybrid Monte Carlo (HMC) with the $\Phi$ hybrid molecular dynamics (MD) algorithm. The $\Phi$ algorithm obtains the force on the gauge field from the derivative of the pseudofermion effective potential $S_{f}=\Phi^{*}[M(U)^{\dagger}M(U)]^{-1}\Phi$ with respect to the vector potential $A$ or $U$. Because the fermion matrix depends on the gauge field, the pseudofermion term contributes the “fermion force”,
@@
A reversible molecular dynamics treatment of a discontinuous effective potential was proposed by [9]. It is based on the analogy with classical particle motion in the presence of a step barrier. Imagine a unit-mass particle moving in two spatial dimensions $(x,y)$, and approaching a potential step, with $V(x,y)=0$, say, for $x<0$ and $V(x,y)=V_{0}$ for $x>0$. The component of momentum in the $y$ direction is unchanged when the particle strikes the barrier, but there is an impulse in the $x$ direction, which causes a discontinuous change in $p_{x}$. If the barrier is too high, the particle “reflects,” $p_{x}\rightarrow-p_{x}$, but if the barrier is low enough, the particle crosses the barrier (“refracts”) while $p_{x}$ changes. The change is given by energy conservation, $\frac{1}{2}p_{x}^{2}=\frac{1}{2}(p_{x}^{\prime})^{2}+V_{0}$.
@@
In practice, the exact step function of the overlap operator is replaced by some approximation, such as a rational functional approximation. To obtain an accurate approximation to the step function, it continues to be necessary to project out low modes of the kernel operator, as in Eq. (61). This introduces another complication for HMC: we have to be able to differentiate the projector. Evaluating the derivative requires determining the change in the projector under small perturbations in $h$. Labeling the projector on an eigenmode of h(r0)h(-r0) with eigenvalue $\lambda$ as $P_{\lambda}=|\lambda\rangle\langle\lambda|$ , first order perturbation theory gives [22]
@@
All this business with reflection and refraction is morally equivalent to the problem of changing topology as seen in simulations with actions which treat topology only approximately (meaning, everything else but overlap fermions). What is different is that here, the barrier between topology changes is sharp and easy to understand. As with other actions, simulations with overlap fermions have a hard time changing topology and the problem gets worse as the lattice spacing decreases. It is very easy, though, to perform simulations in sectors of fixed topology, simply by always reflecting from topological boundaries. These data sets can used in the epsilon regime studies described above, to check predictions for the properties of eigenvalues in sectors of fixed winding number.
The magical properties of overlap fermions allow simulations of an odd number of flavors without rooting. For details, see Ref. [7].
Finally, let me make a few remarks about my own experiences with the overlap as a computational problem. (This is described in more detail in Ref. [6].) I was never able to get anything useful out of an overlap operator whose kernel $d$ or $h$ was a usual Wilson fermion with thin links. It was too expensive to evaluate; $\epsilon_{N}$ needed very large $N$ and very small $c_{2l-1}$’s. The issue, I always thought, was that while the spectrum of the overlap $D$ was a circle, that was not the case for the kernel Wilson action. (An example was shown in Fig. 1.) I always felt that it was asking a lot of the computer to map the crosses in Fig. 1 onto the Ginsparg-Wilson circle. Actions with nearest neighbor and diagonal ($\pm\hat{i}\pm\hat{j}$) couplings produced approximately circular spectra to start with, and it was easier for the overlap operator to circularize them. Smeared gauge connections plus a clover term further smoothed the action and allowed me to use an $\epsilon_{N}$ with small $N$ and a more well-behaved range of eigenvalues of $h$. I used eigenmodes of $h$ to precondition the calculation of $H$ and eigenmodes of $H$ to precondition the calculation of $\hat{D}^{-1}$.
@@ V Final thoughts
As far as I can tell from a literature search, only one lattice group doing large scale simulations used overlap fermions (though there were several groups doing small scale simulations). That was JLQCD, who started carrying out $N_{f}=2$ flavor simulations around 2006, with a $16^{3}\times 32$ volume and quark masses down to a nominal 20 MeV or $m_{q}=m_{strange}/6$ – see Ref. [16]. A year later they were doing $N_{f}=2+1$ flavor simulations. Their overlap program continued until about 2014, at which point they converted to simulating Moebius domain wall fermions [28]. These fermions do not have exact chiral symmetry, but the additive mass renormalization they observed was about 1/2 MeV, which they argued was more than adequate for their needs.
Advances in techniques also rendered many of the theoretical advantages of exact chiral symmetry less important. For example, these days the topological susceptibility can be measured using gradient flow [21] from simulations with nonchiral fermion actions, rather than by counting zero modes. The condensate can be measured from the integrated spectral density of eigenvalues of a (nonchiral) Dirac operator [13]. And of course, most of the “classic problems” from the time of the development of the overlap have been solved: measurements of the condensate or the pseudoscalar decay constant are not so newsworthy any more. Groups have moved on to more complicated observables, where many things have to be balanced against each other – small lattice spacing, large volume, large statistics, multiparticle correlation functions, good chiral symmetry properties, and more.